OpenStudy (anonymous):

First one to answer gets a medal!!! Simplify the decimal number: 04.200 ?

OpenStudy (anonymous):

4.2

OpenStudy (anonymous):

04.200=4.2=42/10=21/5\[=4\frac15\]